Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhGitano Products Schedule of Cost of Goods Manufactured Direct materials: Total raw materials available Raw materials used inCompute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answeAssume that the $38,000 ending balance in Work in Process includes $8,700 of direct materials. Given this assumption, supply
Answer #1
Required 1 :
Predetermined overhead rate = Estimated manufacturing overhead / Estimated direct materials costs = 130200 / 93000 140%
Required 2 :
Beginning raw materials 26000
(+) Purchase of raw materials 136000
Total raw materials available 162000
(-) Ending raw materials 11000
Raw materials used 151000
Overhead applied = Raw materials used * Predetermined overhead rate = 151000 * 140% 211400
Indirect labor 131000
Property taxes 8800
Depreciation of equipment 16000
Maintenance 13000
Insurance 9200
Rent, building 38000
Actual overhead costs 216000
As the overhead applied are less than the actual overhead costs, the overheads are underapplied
Underapplied overhead = Actual overhead costs - Overhead applied = 216000 - 211400 4600
Underapplied Overhead 4600
Required 3 :
Gitano Products
Schedule of cost of goods manufactured
Direct materials :
Beginning raw materials 26000
Purchase of raw materials 136000
   Total raw materials available 162000
Ending raw materials 11000
Raw materials used in production 151000
Direct labor cost 89000
Manufacturing overhead applied 211400
Total manufacturing costs 451400
Beginning work in process 47000
Total cost of work in process 498400
Ending work in process 38000
Cost of goods manufactured 460400
Required 4 :
Cost of goods manufactured 460400
(+) Beginning finished goods 73000
(-) Ending finished goods 55000
Unadjusted cost of goods sold 478400
Required 5 :
Direct materials 8700
Direct labor cost ( 38000 - 8700 - 12180 ) 17120
Manufacturing overhead ( 8700 * 140% ) 12180
Work in process inventory 38000